Q1 Mr. Smith borrowed $189,500 at 3.75% per year compounded monthly. Loan is for 20 years. Compute the monthly payment. Q2 For the same loan in Q1, compute how many months it will take to pay the loan off, if Mr. Smith pays $100 extra to the monthly amount you computed earlierQ3. For the same loan and payment you computed in Q1, compute the amount Mr. Smith owes to the bank immediately after the 89th monthly payment. Q4. Follow up to Q3: Compute the split of interest and principal of the 90th payment.
Answer #1

answer:

Q1

Given

loan amount P=$189,500

Interest rate=3.75% per year

Monthly Interest rate r=3.75%/12=0.3125%

Period N=12*20=240 months

Payment per month A=P*r/(1-(1+r)^-N)

A=189500*0.3125%/(1-(1+0.3125%)^-240)

A=$1123.52

Q2)

Since smith pays $100 extra

So New monthly payment A=100+1123.52=$1223.52

Let n be the number of months to pay off loan

P=A*(1-(1+r)^-n)/r

189500=1223.52*(1-(1+0.3125%)^-n)/0.3125%

-n*ln(1.003125)=ln(0.52)

n=212 months

Q3)

Payment Schedule

Month Balance Loan Amount Payment Per Months Interest contribution Principle Contribution
1 189500.00 1123.52 592.19 531.33
2 188968.67 1124.52 590.53 533.99
3 188434.67 1125.52 588.86 536.66
4 187898.01 1126.52 587.18 539.34
5 187358.67 1127.52 585.50 542.02
6 186816.65 1128.52 583.80 544.72
7 186271.93 1129.52 582.10 547.42
8 185724.51 1130.52 580.39 550.13
9 185174.38 1131.52 578.67 552.85
10 184621.53 1132.52 576.94 555.58
11 184065.95 1133.52 575.21 558.31
12 183507.64 1134.52 573.46 561.06
13 182946.58 1135.52 571.71 563.81
14 182382.77 1136.52 569.95 566.57
15 181816.19 1137.52 568.18 569.34
16 181246.85 1138.52 566.40 572.12
17 180674.73 1139.52 564.61 574.91
18 180099.82 1140.52 562.81 577.71
19 179522.11 1141.52 561.01 580.51
20 178941.59 1142.52 559.19 583.33
21 178358.27 1143.52 557.37 586.15
22 177772.12 1144.52 555.54 588.98
23 177183.13 1145.52 553.70 591.82
24 176591.31 1146.52 551.85 594.67
25 175996.64 1147.52 549.99 597.53
26 175399.11 1148.52 548.12 600.40
27 174798.71 1149.52 546.25 603.27
28 174195.44 1150.52 544.36 606.16
29 173589.28 1151.52 542.47 609.05
30 172980.22 1152.52 540.56 611.96
31 172368.27 1153.52 538.65 614.87
32 171753.40 1154.52 536.73 617.79
33 171135.61 1155.52 534.80 620.72
34 170514.89 1156.52 532.86 623.66
35 169891.23 1157.52 530.91 626.61
36 169264.62 1158.52 528.95 629.57
37 168635.05 1159.52 526.98 632.54
38 168002.51 1160.52 525.01 635.51
39 167367.00 1161.52 523.02 638.50
40 166728.50 1162.52 521.03 641.49
41 166087.01 1163.52 519.02 644.50
42 165442.51 1164.52 517.01 647.51
43 164795.00 1165.52 514.98 650.54
44 164144.46 1166.52 512.95 653.57
45 163490.89 1167.52 510.91 656.61
46 162834.28 1168.52 508.86 659.66
47 162174.62 1169.52 506.80 662.72
48 161511.90 1170.52 504.72 665.80
49 160846.10 1171.52 502.64 668.88
50 160177.22 1172.52 500.55 671.97
51 159505.26 1173.52 498.45 675.07
52 158830.19 1174.52 496.34 678.18
53 158152.02 1175.52 494.23 681.29
54 157470.72 1176.52 492.10 684.42
55 156786.30 1177.52 489.96 687.56
56 156098.73 1178.52 487.81 690.71
57 155408.02 1179.52 485.65 693.87
58 154714.15 1180.52 483.48 697.04
59 154017.11 1181.52 481.30 700.22
60 153316.90 1182.52 479.12 703.40
61 152613.49 1183.52 476.92 706.60
62 151906.89 1184.52 474.71 709.81
63 151197.08 1185.52 472.49 713.03
64 150484.05 1186.52 470.26 716.26
65 149767.79 1187.52 468.02 719.50
66 149048.30 1188.52 465.78 722.74
67 148325.55 1189.52 463.52 726.00
68 147599.55 1190.52 461.25 729.27
69 146870.28 1191.52 458.97 732.55
70 146137.73 1192.52 456.68 735.84
71 145401.89 1193.52 454.38 739.14
72 144662.75 1194.52 452.07 742.45
73 143920.30 1195.52 449.75 745.77
74 143174.53 1196.52 447.42 749.10
75 142425.43 1197.52 445.08 752.44
76 141672.99 1198.52 442.73 755.79
77 140917.20 1199.52 440.37 759.15
78 140158.05 1200.52 437.99 762.53
79 139395.52 1201.52 435.61 765.91
80 138629.61 1202.52 433.22 769.30
81 137860.31 1203.52 430.81 772.71
82 137087.60 1204.52 428.40 776.12
83 136311.48 1205.52 425.97 779.55
84 135531.93 1206.52 423.54 782.98
85 134748.95 1207.52 421.09 786.43
86 133962.52 1208.52 418.63 789.89
87 133172.64 1209.52 416.16 793.36
88 132379.28 1210.52 413.69 796.83
89 131582.45 1211.52 411.20 800.32
90 130782.12 1212.52 408.69 803.83

After 89th payment Mr. Smith owes $130782.12 to bank.

Q4)

for 90th payment

Interest =$408.69

Principle=$803.83